Tickets to a school dance cost $3, and the projected attendance is 325 persons. It is further projected that for every 10¢ increase in ticket price, the average attendance will decrease by 5. At what ticket price will the receipts from the dance be $1,125?
Answer by lwsshak3(11628)

Tickets to a school dance cost $3, and the projected attendance is 325 persons. It is further projected that for every 10¢ increase in ticket price, the average attendance will decrease by 5. At what ticket price will the receipts from the dance be $1,125?
***
let x=number of 10ç increases in ticket price
Receipts=ticket price*attendance
R(x)=(3+.10x)(325-5x)=1125
R(x)=975+32.5x-15x-.5x^2=1125
-.5x^2+17.5x-150=0
-5x^2+175x-1500=0
-x^2+35x-300=0
(-x+15)(x-20)=0
x=15*10ç=1.50
ticket price=3.00+1.50=$3.50
or
x=20*10ç=2.00
ticket price=3.00+2.00=$5.00
At what ticket price will the receipts from the dance be $1,125? $3.50 or $5.00